OpenStudy (anonymous):

5x+7y=31 2x+4y=16 solve the given system of equations by the addition method?

OpenStudy (anonymous):

multiply first equation by 2 and second by (-5) and add together. what would you get?

OpenStudy (anonymous):

10x+14y=62 -10x-20y=-80 ------------- (+)

OpenStudy (anonymous):

-6y=-18 y=....?

OpenStudy (anonymous):

y=3 substitute value of y=3 in any given equation - I'll do it with the first one: 5x+7*3=31 5x=31-21 x=10/5 x=2 So the final answer would be: x=2 and y=3 Got to go. Good Luck!